Circular No. 6616

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997bq IN NGC 3147 Stephen Laurie, Church Stretton, Shropshire, U.K., reports his discov-

ery of a supernova (mag ∼ 16.1) on CCD images taken with a 0.25-m reflector during Apr. 7.91–8.00 UT. SN 1997bq is located at α = 10^{h}17^{m}04^{s}.0, δ = +73^{o}23′03′′ (equinox 2000.0), which is 1′ southeast of the galaxy’s nucleus, embedded in the fainter outer arms of the galaxy. A nearby star (mag 18) has position end figures 07^{s}.1, 10′′. No star is visible at the position of the supernova on images taken on Feb. 27 and Mar. 24.91 (limiting mag 17.0–17.5), nor is anything present on the Digital Sky Survey.

P. Challis, Harvard-Smithsonian Center for Astrophysics, reports that a spectrum obtained at the Multiple Mirror Telescope shows this to be a type-Ia supernova, 1 or 2 days before maximum.

GRB 970402 P. J. Groot and T. J. Galama, University of Amsterdam (UoA); J. van

Paradijs, UoA and University of Alabama in Huntsville; C. Kouveliotou, Universities Space Research Association, at Marshall Space Flight Center, NASA; R. Strom, Netherlands Foundation for Research in Astromony; M. Brown and R. Webster, University of Melbourne; L. Van Zyl and B. Warner, University of Cape Town; J. Menzies, South African Astronomical Observatory; J. in ’t Zand and J. Heise, Space Research Organization of the Netherlands, Utrecht; E. Costa, M. Feroci, and L. Piro, Istituto di Astrofisica Spaziale, CNR; F. Frontera and G. Zavattini, Università di Ferrará; and L. Nicastro and E. Palazzi, Istituto Tecnologia e Studie Radiazione Extraterrestri, CNR, report: “We obtained R-band images of the 3′-radius error box of GRB 970402 (IAUC 6610) on Apr. 3.7 UT (2400 s, limiting mag R = 21.5), and on Apr. 4.8 (3000 s, limiting mag 21.0, due to high winds) using the 1.0-m telescope at Siding Spring. We also obtained an R-band mosaic of five pointings covering 35 percent of the error box, with integration times of 45 min per pointing (limiting mag R = 21.5), with the South African Astronomical Observatory’s 1-m telescope on Apr. 4.0. A comparison among these frames shows no obvious increase or decrease in brightness in any object by more than 0.3 mag (2σ), to a limit of R = 21.0. Comparison with the Digital Sky Survey shows no brightening of sources to R = 19.”

1997 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green